OpenStudy (anonymous):

Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Can you show me a problem that has a binomial that is a factor of the polynomial being divided, and another problem that has a binomial that is not a factor of the polynomial being divided.

OpenStudy (anonymous):

Somebody please help me.

OpenStudy (anonymous):

Can't really tell until the division is done

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

You have to divide the polynomial by the binomial first.

OpenStudy (anonymous):

I mean in general there's no rule/trick that helps you detect if a binomial is a factor of a polynomial, you have to do the dividing work yourself to see it.

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Yes, in general.

OpenStudy (anonymous):

This is part of a project I'm doing. I don't think your answer is helping.

OpenStudy (anonymous):

What is that for?

OpenStudy (anonymous):

Yes, I think so. But if you come up with such a rule to see if a binomial is a factor of polynomial without actually dividing it, feel free to publish your work here!

OpenStudy (anonymous):

Can you at least help me with this: x^16 - 1/ x - 1. Simplify this by factoring and long division. @drawar

OpenStudy (anonymous):

Well, this is a case when you have to tell immediately that x-1 is a factor of x^16 -1, and also you should write the factorization in a flash: \[=(x^{8}+1)(x^{4}+1)(x^{2}+1)(x+1)\]

OpenStudy (anonymous):

What's the next step?

OpenStudy (anonymous):

What next step? x^16/x-1 = that stuff I wrote over there

OpenStudy (anonymous):

But don't you have to simplify this: (x^8+1)(x^4+1)(x^2+1)(x+1) ?

OpenStudy (anonymous):

What method did you use for this problem?

OpenStudy (anonymous):

well, if you want, you can use complex numbers to factor out all these high degree binomials into binomials of degree 1

OpenStudy (anonymous):

There's no magic here, just repeated uses of the famous identity a^2-b^2 = (a-b)(a+b)

OpenStudy (anonymous):

I didn't get what you were saying in the first comment. But I'm only a little familiar with what you said in the second comment.

OpenStudy (anonymous):

take x^2+1 for example, there's no real roots to the eq. x^2+1=0, you can factor it out using real numbers anymore. So if needed, you can make use of the complex number i, whose square equals -1, and write x^2+1 as (x+i)(x-i)

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Which is easier to use for this problem: x^16 - 1/ x - 1, factoring or long division?

OpenStudy (anonymous):

factoring

OpenStudy (anonymous):

Is what you showed you using factoring?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The last part which was (x + 1) shouldn't the x be x^2, because if you add the exponents that way you get x^16. When you typed x, I added up the exponents and got 15 instead of 16. (X^15 instead of x^16).

OpenStudy (anonymous):

What I wrote is the result after simplifying the expression x^16-1/x-1 using factoring. If you want to get x^16-1, you have to multiply it by x-1, makes sense?

OpenStudy (anonymous):

Can you show me how you did the multiplication? So I can see how you worked It out.

OpenStudy (anonymous):

@drawar

OpenStudy (anonymous):

just do from left to right: (x-1)(x+1)(x^2+1)(x^4+1)(x^8+1)=(x^2-1)(x^2+1)(x^4+1)(x^8+1)=(x^4-1)(x^4+1)(x^8+1)=(x^8-1)(x^8+1)=x^16-1

OpenStudy (anonymous):

Thanks. And this is using the factoring method, right?

OpenStudy (anonymous):

Yes, because when you look at x^16-1, you immediately know what its factorization looks like. Long division would look messy in this case.

OpenStudy (anonymous):

Thanks